Angles A,B and C of a trangle ABC are in A.P. if b/c=3^1/2/2^1/2 then angle A is equal to

  1. Angles A,B and C of a trangle ABC are in A.P. if b/c=3^1/2/2^1/2 then angle A is equal to

Grade:11

1 Answers

Aditya Gupta
5 years ago
hi aditya, im aditya too.
let the angles be x-a, x, x+a
adding shud give 180
so  x-a+ x+ x+a=180
or x=60
so B is 60 deg
by sin rule we have sinB/sinC=b/c=√(3/2)
but sinB=√3/2
or sinC=1/√2
so C=45
so A=180-(60+45)
A=75degree